# Cross-correlating IceCube neutrinos with a large set of galaxy samples around redshift $z\sim 1$

### v1

@@ IV.3 Angular power spectra and covariance matricies
An observed field $\tilde{X}$ with incomplete sky coverage can be modelled as the product of the true underlying field $X$ with a mask $w$. Often the the mask is treated simply as a binary map where $w=1$ for pixels on the sky that have been observed and $w=0$ where they have not, but in general $w$ can be designed to optimally weight different regions of the sky based on varying noise levels. The pseudo-$C_{\ell}$ for two fields $X$ and $Y$ is then defined as
@@
where $M_{\ell\ell^{\prime}}$ is the mode-coupling matrix that can be computed directly from the power spectra of the masks. Eq. 12 is not directly invertible due to the loss of information on the masked sky, so the pseudo-$C_{\ell}$ is binned into bandpowers by assuming that it is piecewise constant in a number of discrete bins and a “binned” mode-copling matrix is computed that can be inverted to obtain an estimate of the true power spectrum. If one or both of the fields are smoothed by a PSF, then we divide the pseudo-$C_{\ell}$ by the corresponding beam window function before inverting the mode coupling matrix.
@@
We binned the power-spectra using linear bins of width $\Delta\ell=50$ in the range $2\leq\ell<3N_{\text{side}}$. While we calculated the cross-spectra all the way out to $\ell=1536$ to avoid power leakage, we discard all bandpowers above $\ell=350$ for the analysis in order to avoid systematic errors from mis-characterizing the IceCube PSF at small scales and from the breakdown of the linear bias assumption at non-linear scales. We additionally discard the first bandpower at $\ell<50$ in order to avoid potential large-scale systematics in the galaxy maps and the neutrino mean map.
@@ V.1 Model fit
where $\mathbf{d}$ is the data vector containing all of the meeasured galaxy-neutrino cross-spectra between $35050<\ell<$ 350 and $C$ is the full covariance matrix. Here we have ignored any uncertainty in the galaxy or CIB models since the data is completely dominated by noise from the neutrinos.
@@ V.2 Energy binning
The quantity $f_{\text{astro}}$ for track-like events is expected to depend very strongly on energy, due to the softer energy spectrum of atmospheric neutrinos compared to the astrophysical ones. Here we divide the neutrino events into three equal-size logarithmic energy bins between $10^{3}$ GeV and $10^{6}$ GeV. These bins have roughly 390,000, 9,700, and 218 events respectively. Based on Ref. [2], we expect that $f_{\text{astro}}$ for each bin to be on the order of 0.006, 0.06, and 0.45 respectively. For each energy bin, we construct overdensity maps, estimate the effective beam, and compute cross-spectra with the galaxy samples in the exact same manner as described for the entire neutrino sample.
We again use the three assumed models for the neutrino redshift distribution and plot the resulting constraints on $b_{\nu}f_{\text{astro}}$ for the three energy bins in the top row of Fig. 8 and compare them to the expected values of $f_{\text{astro}}$ from energy spectrum measurements. We find that in the simple “peak at $z\sim 1$” model, the $2\sigma$ significant correlation between neutrinos and LSS is driven entirely by the lowest energy bin, potentially in slight conflict with expectations based on the energy spectrum measurement of $f_{\text{astro}}$. On the other hand, in the flux-weighted models we get potential detections coming from the high energy events. In the “no evolution” (“following SFR”) model, the lowest and highest energy bins are significant at $1\sigma$ and $1.9\sigma$ ($1.5\sigma$ and $1.6\sigma$) respectively. All of these significances are low and the $\Delta\chi^{2}$ values for each of the models are not different enough to prefer one of them over the other. At this point, it is unclear whether we have found a real correlation between neutrinos and LSS or simply a noise fluctuation.
@@ V.3 Forecasts
We consider three possible scenarios for future neutrino detectors. The first scenario assumes the same number of detected events as the currrent 10-year dataset, but with 3 times better angular uncertainties. The second scenario, “Gen2”, assumes 10 times as many events with 3 times better angular uncertainties (roughly in line with what is expected from IceCube-Gen2 [52]). Finally, the third scenario, “Gen2 ideal”, assumes 10 times as many events with essentially no angular uncertainties. We then calculate the expected errors on the measurements of $b_{\nu}f_{\text{astro}}$, $\sigma(b_{\nu}f_{\text{astro}})$, assuming ideal measurements of the cross-correlations out to $\ell_{\text{max}}=750$. As a check, we confirm that a forecast of the expected constraining power of current data matches the constraints we got in the analysis of actual data.
We plot the resulting forecasts of $\sigma(b_{\nu}f_{\text{astro}})$ compared with the errors obtained in the current analysis in the bottom row of Fig. 8. As expected, in the noise-dominated regime, we find that $\sigma(\nu f_{\text{astro}})$ is independent of $b_{\nu}f_{\text{astro}}$ and scales linearly with $1/\sqrt{N_{\nu}}$, where $N_{\nu}$ is the total number of neutrinos in the sample. We find that a Gen2-like detector should provide measurements of the galaxy-neutrino cross-spectra with roughly 5 to 6 times better SNR than the current measurements, which should be sufficient to probe interesting values of $b_{\nu}f_{\text{astro}}$. In addition, future data would provide significatly better power to discriminate between different models of the neutrino redshift distribution.
@@ VI Discussion
In our analysis we are not able to meaningfully constrain the redshift distribution of astrophysical neutrinos. Physical models of neutrino production in astrophysical sources could help put theoretical priors on the neutrino source redshift distribution. This, in turn, will allow better constraints on $b_{\nu}$ as well as on the shape of the neutrino redshift distribution $p_{\nu}(z)$. Ultimately, this will allow us to pin down the astrophysical populations that build up the diffuse astrophysical neutrino flux.
While the current analysis of neutrino-galaxy cross-correlations is severly limited by the quality of the neutrino data, there are some improvements that could be done on the galaxy side as well. Given that the flux-weighted neutrino distributions are heavily skewed to low-redshift (Fig. 7), detailed analysis using low-redshift galaxy samples, especially the DESI Bright Galaxy Sample [53], will be important to discriminate between these models. At low redshifts, though, more careful modelling than what we have done will be necessary in order to account for the more non-linear and non-Gaussian galaxy fields. While we have used a few low-redshift samples in this work, the majority of our galaxy samples are at significantly higher redshifts where the simple linear bias model should be adequate on the scales we have considered. In the future it will also be necessary to do a full combined fit for the galaxy-neutrino cross-correlations, taking into account the uncertainties in the galaxy data.
@@ Acknowledgements.
We thank Alex Krolewski for sharing the unWISE galaxy maps and providing comments on our manuscript. We also thank Abhishek Maniyar and Ziang Yan for guidance with the CIB model. This work was partially supported by the Center for AstroPhysical Surveys (CAPS) at the National Center for Supercomputing Applications (NCSA), University of Illinois Urbana-Champaign. A.O. acknowledges support from the Illinois Center for Advanced Studies of the Universe (ICASU)/CAPS/NCSA Graduate Fellowship. This work made use of the Illinois Campus Cluster, a computing resource that is operated by the Illinois Campus Cluster Program (ICCP) in conjunction with the National Center for Supercomputing Applications (NCSA) and which is supported by funds from the University of Illinois at Urbana-Champaign. This research has made use of NASA’s Astrophysics Data System and adstex (https://github.com/yymao/adstex). We also acknowledge the open-source software that was invaluable in this research, including: Python, NumPy [54], SciPy [55], and matplotlib [56].

### v2

@@ IV.3 Angular power spectra and covariance matrices
An observed field $\tilde{X}$ with incomplete sky coverage can be modelled as the product of the true underlying field $X$ with a mask $w$. Often the mask is treated simply as a binary map where $w=1$ for pixels on the sky that have been observed and $w=0$ where they have not, but in general $w$ can be designed to optimally weight different regions of the sky based on varying noise levels. The pseudo-$C_{\ell}$ for two fields $X$ and $Y$ is then defined as
@@
where $M_{\ell\ell^{\prime}}$ is the mode-coupling matrix that can be computed directly from the power spectra of the masks. Eq. 26 is not directly invertible due to the loss of information on the masked sky, so the pseudo-$C_{\ell}$ is binned into bandpowers by assuming that it is piecewise constant in a number of discrete bins and a “binned” mode-coupling matrix is computed that can be inverted to obtain an estimate of the true power spectrum. If one or both of the fields are smoothed by a PSF, then we divide the pseudo-$C_{\ell}$ by the corresponding beam window function before inverting the mode coupling matrix. We do not include the effect of finite map resolution at this stage, but instead include the pixel window function in the model for the cross-correlations, though we note that this effect is completely negligible ($<1$% change in our final results) at $N_{\text{side}}=512$ due to the aggressive smoothing introduced by the IceCube effective beam.
@@
We binned the power-spectra using linear bins of width $\Delta\ell=50$ in the range $2\leq\ell<3N_{\text{side}}$. While we calculated the cross-spectra all the way out to $\ell=1536$ to avoid power leakage, we discard all bandpowers above $\ell=350$ for the analysis in order to avoid systematic errors from mis-characterizing the IceCube PSF at small scales and from the breakdown of the linear bias assumption at non-linear scales. At the same time, all of the signal is concentrated at low $\ell$ due to the aggressive smoothing of the neutrino map, so we see almost no change in our final results when significantly increasing this upper scale cut. We additionally discard the first bandpower at $\ell<50$. This is a relatively conservative choice driven by several considerations. First, this reduces any systematic effects from large spatial variations in the effective beam. As seen in Fig. 4, our fiducial effective beam (using the auto-spectra of the neutrino maps) begins to diverge from the analytic estimate (assuming no spatial variation) below $\ell=50$. This $-\ell$ cut is also used when estimating the auto-spectra of the galaxy samples in order to avoid potential large-scale systematics in the galaxy maps, and allows us to safely use the Limber approximation (Eq. 2) when modelling all of the power spectra. Finally, we note that in our analysis, there are negligible changes in our final results whether we include or discard this first bandpower. In future analyses with higher sensitivity, it will be necessary to more rigorously test all of these assumptions that go into choosing scale cuts.
@@ V.1 Model fit
where $\mathbf{d}$ is the data vector containing all of the measured galaxy-neutrino cross-spectra between $35050<\ell<$ 350 and $C$ is the full covariance matrix. To confirm that this assumption of Gaussianity is appropriate, we have run a Kolmogorov-Smirnov test on the $b_{\nu}f_{\text{astro}}$ values resulting from 25 random neutrino map simulations, and for all $p_{\nu}(z)$ models, find $p$-values $>0.4$. We found similarly high $p$-values for the energy-binned neutrino maps that are presented in the following section. Thus, when computing error bars or upper limits on $b_{\nu}f_{\text{astro}}$, we simply assume Gaussian distributions in all cases. We have also ignored any uncertainty in the galaxy or CIB models since the data is completely dominated by noise from the neutrinos.
@@ V.2 Energy binning
The quantity $f_{\text{astro}}$ for track-like events is expected to depend very strongly on energy, due to the softer energy spectrum of atmospheric neutrinos compared to the astrophysical ones. Thus, to increase our sensitivity, we divide the neutrino events into three equal-size logarithmic energy proxy bins between $10^{3}$ GeV and $10^{6}$ GeV. These bins have roughly 390,000, 9,700, and 218 events respectively. We again emphasize that this energy proxy does not directly correspond to the true neutrino energy (see Ref [19]), but higher energy proxies do generally correlate with higher true neutrino energies. Based on Ref. [2], which measures the energy spectrum of a somewhat similar event selection to ours (track-like events in the northern sky), we expect that $f_{\text{astro}}$ for each bin to be on the order of 0.006, 0.06, and 0.45 respectively.
For each energy bin, we construct overdensity maps, estimate the effective beam, and compute cross-spectra with the galaxy samples in the exact same manner as described for the entire neutrino sample.
We again use the three assumed models for the neutrino redshift distribution and plot the resulting constraints on $b_{\nu}f_{\text{astro}}$ for the three energy bins in the top row of Fig. 9 and compare them to the expected values of $f_{\text{astro}}$ from energy spectrum measurements. We find that in the simple “peak at $z\sim 1$” model, the $2\sigma$ significant correlation between neutrinos and LSS is driven entirely by the lowest energy bin, potentially in slight conflict with expectations based on the energy spectrum measurement of $f_{\text{astro}}$. On the other hand, in the flux-weighted models we get potential detections coming from the high energy events. In the “no evolution” (“following SFR”) model, the lowest and highest energy bins are significant at $1\sigma$ and $1.9\sigma$ ($1.5\sigma$ and $1.6\sigma$) respectively. All of these significance values are low and the $\Delta\chi^{2}$ values for each of the models are not different enough to prefer one of them over the other. At this point, it is unclear whether we have found a real correlation between neutrinos and LSS or simply a noise fluctuation.
@@ V.3 Forecasts
We consider three possible scenarios for future neutrino detectors. The first scenario assumes the same number of detected events as the current 10-year dataset, but with 3 times better angular uncertainties. The second scenario, “Gen2”, assumes 10 times as many events with 3 times better angular uncertainties (roughly in line with what is expected from IceCube-Gen2 [52]). Finally, the third scenario, “Gen2 ideal”, assumes 10 times as many events with essentially no angular uncertainties. We then calculate the expected errors on the measurements of $b_{\nu}f_{\text{astro}}$, $\sigma(b_{\nu}f_{\text{astro}})$, assuming ideal measurements of the cross-correlations out to $\ell_{\text{max}}=750$. As a check, we confirm that a forecast of the expected constraining power of current data matches the constraints we got in the analysis of actual data.
We plot the resulting forecasts of $\sigma(b_{\nu}f_{\text{astro}})$ compared with the errors obtained in the current analysis in the bottom row of Fig. 9. As expected, in the noise-dominated regime, we find that $\sigma(b_{\nu}f_{\text{astro}})$ is independent of $b_{\nu}f_{\text{astro}}$ and scales linearly with $1/\sqrt{N_{\nu}}$, where $N_{\nu}$ is the total number of neutrinos in the sample. We find that a Gen2-like detector should provide measurements of the galaxy-neutrino cross-spectra with roughly 5 to 6 times better SNR than the current measurements, which should be sufficient to probe interesting values of $b_{\nu}f_{\text{astro}}$. In addition, future data would provide significantly better power to discriminate between different models of the neutrino redshift distribution.
@@ VI Discussion
In our analysis we are not able to meaningfully constrain the redshift distribution of astrophysical neutrinos. Physical models of neutrino production in astrophysical sources could help put theoretical priors on the neutrino source redshift distribution. This, in turn, will allow better constraints on $b_{\nu}$ as well as on the shape of the neutrino redshift distribution $p_{\nu}(z)$. Ultimately, this will allow us to pin down the astrophysical populations that build up the diffuse astrophysical neutrino flux. This approach is also very similar to that taken by Ref. [53], where the authors drew the connection between neutrino sky anisotropies and a possible cutoff in the energy spectrum of ultrahigh energy neutrinos.
While the current analysis of neutrino-galaxy cross-correlations is severely limited by the quality of the neutrino data, there are some improvements that could be done on the galaxy side as well. Given that the flux-weighted neutrino distributions are heavily skewed to low-redshift (Fig. 8), detailed analysis using low-redshift galaxy samples, especially the DESI Bright Galaxy Sample [54], will be important to discriminate between these models. At low redshifts, though, more careful modelling than what we have done will be necessary in order to account for the more non-linear and non-Gaussian galaxy fields. While we have used a few low-redshift samples in this work, the majority of our galaxy samples are at significantly higher redshifts where the simple linear bias model should be adequate on the scales we have considered. In the future it will also be necessary to do a full combined fit for the galaxy-neutrino cross-correlations, taking into account the uncertainties in the galaxy data.
@@ Acknowledgements.
We thank Alex Krolewski for sharing the unWISE galaxy maps and providing comments on our manuscript. We also thank Abhishek Maniyar and Ziang Yan for guidance with the CIB model. We thank the anonymous referee for providing comments that helped improve this paper. This work was partially supported by the Center for AstroPhysical Surveys (CAPS) at the National Center for Supercomputing Applications (NCSA), University of Illinois Urbana-Champaign. A.O. acknowledges support from the Illinois Center for Advanced Studies of the Universe (ICASU)/CAPS/NCSA Graduate Fellowship. This work made use of the Illinois Campus Cluster, a computing resource that is operated by the Illinois Campus Cluster Program (ICCP) in conjunction with the National Center for Supercomputing Applications (NCSA) and which is supported by funds from the University of Illinois at Urbana-Champaign. This research has made use of NASA’s Astrophysics Data System and adstex (https://github.com/yymao/adstex). We also acknowledge the open-source software that was invaluable in this research, including: Python, NumPy [55], SciPy [56], and matplotlib [57].
Appendix A Theory curves for cross-spectra
In Fig. 10 we show all of the theory curves for the neutrino-galaxy cross-spectra computed for each of the model neutrino redshift distributions that we have tested. These curves are calculated for $b_{\nu}f_{\text{astro}}=1$ so that they can be used as model templates in Eqs. 29 and 30.
Figure 10: Theory curves (un-binned) for the neutrino-galaxy cross-spectra $C_{\ell}^{g\nu}$. Each column corresponds to a different model for the neutrino redshift distribution (as given in Section V.1), while each row corresponds to a different set of galaxy samples as given in the legends. These curves are all computed assuming $b_{\nu}f_{\text{astro}}=1$ and linear bias for the galaxies as given in Table 1. Scanning across the columns, it is easy to see how changing the model for the neutrino redshift distribution significantly impacts the relative amplitudes of the cross-spectra.